Suppose we want to determine the (binomial) probability (p) of getting 8 heads in 10 flips of a 2-sided coin. Using Table 2 in Appendix B of the text, what values of n, x and p would we use to look up this probability, and what would be the probability?
Answer by stanbon(75887)

Suppose we want to determine the (binomial) probability (p) of getting 8 heads in 10 flips of a 2-sided coin.
------
Binomial Problem with n = 10; p(head) = 1/2 ; x = 8
-----
P(x = 8) = 10C8(1/2)^8(1/2)^4 = 0.0439
------------------------
Cheers,
Stan H.
===========